# A LLM-Powered Automatic Grading Framework with Human-Level Guidelines Optimization

### v1

@@ 4.2.1. Agent Configurations
The Grader focuses on mapping $x_{i}$ to $s_{j}$ based on the given $G$. In GradeOpt, we leverage the exceptional instruction-following capability of LLMs by using a prompt to instruct LLMs to simulate the grading process of human graders. To fully exploit the potential of LLMs, we incorporate the prompt engineering strategy Chain-of-Thought (Wei et al., 2023). This encourages LLMs to provide both judgment and intermediate reasoning steps in their outputs. With such design, the Grader becomes better aligned with the human-like grading process. Meanwhile, the intermediate reasoning steps provide support for the Reflector to discover the potential improvements to the given guideline. The prompt for the Grader agent is shown in Figure 3.
Task: In this task, you perform the task of assessing teachers’ knowledge of students’ mathematical thinking by grading teacher’s response to a math teaching question. Question: ¡QUESTION¿ Key Concept: ¡KEY CONCEPT¿ Rubrics: ¡RUBRICS¿ Rules: ¡RULES¿ Output format ¡SCORE¿ Reasoning: ¡REASONING¿ Output Rules 1. Repalce ”¡SCORE¿” with ONLY ONE INTEGER from 0, 1, or 2. 2. Replace ”¡REASONING¿” with your reasoning for score assignment. 3. **You MUST strictly obey this output format.** The first token in your response MUST be an integer score. Let’s think step by step!
Figure 3. An illustration of the prompt to LLM-based Grader.
@@
The role of Reflector is to propose ways to improve the current guideline $G_{t-1}$ by reflecting over the error samples returned by Grader. To be specific, we design a two-step instruction prompt for LLMs to achieve this goal. In the first step, LLM is instructed to analyze the individual and shared failure reasons for a set of error samples. Then, in the second step, we ask LLMs to propose suggestions that can help resolve those issues. In general, the two-step improving process is analogous to the gradient descent algorithm used by parameter optimization for machine learning algorithm (Ruder, 2017). In our case, the guideline $G$ serves as the parameter of Grader and identifying the error reason is similar to the ”gradient”. Finally, proposing improving suggestions based on discovered reasons is similar to making a descent down the ”gradient” and thus optimizing $G_{t-1}$. The prompt for the Reflector agent is shown in Figure 4.
CONTEXT You are ReflectorGPT, a helpful AI agent capable of reflecting on [ADAPTATION RUBRICS] that is used by a zero-shot classifier for a grading task. Your task is to reflect and give reasons for why [ADAPTATION RUBRICS] have gotten the given examples in [FAILED EXAMPLES] wrong. OBJECTIVE I’m trying to write a zero-shot classifier for a grading task. The prompt contains two components: 1. [QUESTION STEM], [KEY CONCEPT] and [SCORING RUBRICS] (these three are given by experts and should not be modified); 2. [ADAPTATION RUBRICS] (your task to modify).
Important Steps For Devising Rules: Read [FAILED EXAMPLES]. For each one of the errors, carefully perform the following steps: - Step 1: Clearly explain why the classifier made the mistakes, and provide detailed, explanative analyses of why this teacher response should not be interpreted in that wrong way. - Step 2: Devise or modify [ADAPTATION RUBRICS] for each mistake to help classifier effectively avoid the mistake and classify the teacher response into the correct category (label). Make sure the devised rule is explanative, straightforward, detailed, concise, and in 1 to 3 sentences.
Question Stem: ¡QUESTION STEM¿ Key Concept: ¡KEY CONCEPT¿ Scoring Rubrics: ¡SCORING RUBRICS¿ Adaptation Rubrics: ¡ADAPTATION RUBRICS
But [ADAPTATION RUBRICS] gets the following examples wrong: Failed Examples: ¡ERRORS¿
Give reasons for why [ADAPTATION RUBRICS] could have gotten each and all of these examples wrong.
OUTPUT FORMAT Wrap your response with ¡START¿ and ¡END¿.
START ANALYSIS Let’s think step by step!
Figure 4. An illlustration of the prompt to LLM-based Reflector.
@@
CONTEXT I’m trying to write a zero-shot classifier for a grading task. You are RefinerGPT, a helpful AI agent capable of refining [ADAPTATION RUBRICS] to be used by the zero-shot classifier for giving the most accurate predictions.
OBJECTIVE [ADAPTATION RUBRICS] must contain patterns learned from failed examples, explaining why the predicted score is wrong comparing to the correct label. You should generate new rules based on [ERROR FEEDBACKS] so that the classifier can effectively avoid the mistakes by using your new rules. Your new rules must strictly abide by [SCORING RUBRICS]. Note: Do NOT refer to ”Example 1” ”Example 2” etc. without giving details in your generated rules - when addressing these examples you must cite details to give detailed explanations.
Question Stem: ¡QUESTION STEM¿ Key Concept: ¡KEY CONCEPT¿ Scoring Rubrics: ¡SCORING RUBRICS¿ Adaptation Rubrics: ¡ADAPTATION RUBRICS¿
But [ADAPTATION RUBRICS] have gotten several examples wrong, with the reasons of the problems examined as follows: Failed Examples: ¡ERRORS¿ Error Feedbacks: ¡ERROR FEEDBACKS¿
Based on the above information, I wrote one different improved set of rules in replacement of [ADAPTATION RUBRICS] for instructing the zero-shot classifier to learn patterns from examples for avoiding such errors.
Important: The set of rules must strictly abide by [SCORING RUBRICS] and must clearly use patterns/details from examples to clearly illustrate and explain. However, you must note that [ADAPTATION RUBRICS] are different from [SCORING RUBRICS] or [KEY CONCEPT] - [ADAPTATION RUBRICS] must clearly explain how to correctly classify based on the patterns from failed examples, as learned from [ERROR FEEDBACKS].
The improved set of rules is wrapped with ¡START¿ and ¡END¿. The improved set of new rules are:
Let’s think step by step!
Figure 5. An illustration of the prompt to LLM-based Refiner.
4.2.2. Iterative Optimization Designs
@@
To accelerate the optimization process and encourage a wider exploration of all possible combinations of error samples in $b_{out}$, we integrate the beam searching strategy (Freitag and Al-Onaizan, 2017) within both inner and outer loops. The algorithm of the nested iteration is shown in Algorithm  1. To be specific, in the $w$-th inner loop of the $t$-th outer iteration, GradeOpt accepts guidelines beam $G_{t,w-1}=\{g_{t,w-1}^{(k)}\mid 1\leq k\leq K\}$ from $(w-1)$-th inner iteration instead of a single guideline for refining (line 5). Then, during the inner iteration, each $g_{t,w-1}^{(k)}$ will be sent for reflection and refinement with $L$ independently sampled inner batches $b_{in}$ in a parallel manner (line 9). After all refined guidelines for the $w$-th inner loop are finished, $G_{t,w}=\{g_{t,w}^{(l,k)}\mid 1\leq l\leq L,1\leq k\leq K\}$, each new guideline $g_{t,w}^{(l,k)}$ will be tested over a hold-out validation set $\mathcal{D}_{val}$ (line 14). Meanwhile, the top-$K$ performing guidelines will be kept as $G_{t,w}$ and passed to the $(w+1)$-th inner loop. Finally, the beam output of the last iteration of inner loop $G_{t,W}$ will be sent to the $(t+1)$-th outer iteration (line 4).
@@
While this procedure helps increase the accuracy and reliability, blindly increasing the iteration could lead to over-fitting and higher computational overheads(Juneja et al., 2024). This is particularly true for smaller datasets. To help address these challenges, we introduce an early-stopping criteria. Specifically, during the selection for top-K performed $G_{t,w}$ in the $w$-th inner loop, we record the performance metric $m_{w}$ of the best performed guideline. Then, in the next $(w+1)$-th inner iteration, we check if $m_{w+1}$ is improved. If $m_{w}$ stops improving for two consecutive inner iterations, it indicates that the current guideline is facing risks to be over-fitted, thus following inner iterations are skipped. Similarly, during the $t$-th outer iterations, if its inner iteration is terminated due to the early-stopping and $(t-1)$-th outer iteration’s inner iteration is also terminated by early-stopping, the following outer iterations will also be skipped.
@@
1 Initialize $G_{0,W}=\{g^{(k)}_{0,W}\}=\{\mathcal{G}\}$;
2 for $t\leftarrow 1$ to $T$ do
3       $b_{out}\leftarrow$ sample an outer iteration batch from $\mathcal{D}_{train}$ ;
4       Initialize $G_{t,0}$ = $G_{t-1,W}$ ;
5       for $w\leftarrow 1$ to $W$ do
6
7            for $k\leftarrow 1$ to $K$ do
8                   $\hat{y}_{out}\leftarrow$ generate grading results for $b_{out}$ by Grader with guideline $g^{(k)}_{t,w}$ ;
9                   $e_{t,k}\leftarrow$ find error graded samples from $b_{out}$ caused by guideline $g^{(k)}_{t,w}$ ;
10                   do $l\leftarrow 1$ to $L$ in parallel
11                         $b_{in}^{(l)}\leftarrow$ randomly sample an inner batch from $e_{t,k}$ ;
12                         $g_{t,w}^{(k,l)}\leftarrow$ generate optimized guideline with inputting $b_{in}^{(l)}$ and $g_{t,w-1}^{(k)}$ to Reflector and Refiner ;
13
14                   end
15
16             end for
17            $G_{t,w}=\{g_{t,w}^{(k)}\mid 1\leq k\leq K\}\leftarrow$ select top-$K$ performed $g_{t,w}^{(k,l)}$ based on grading performance over $\mathcal{D}_{val}$ ;
18
19       end for
20
21 end for
@@ 11                         $b_{in}^{(l)}\leftarrow$ randomly sample an inner batch from $e_{t,k}$ ;
Batch Sampling Strategies
Using self-reflective approaches of LLMs to refine grading guidelines requires the exposure of similar errors in consecutive optimization iterations due to LLMs’ lack of ability in generating appropriate modifications with one attempt (Ma et al., 2024). This is especially true for complicated cases involving nuanced differences between score categories. However, the randomness of batch sampling in the outer loop fails to guarantee this pre-requisite, which limits the performance of GradeOpt. To solve this, we develop a novel sampling strategy, which leverages the misconfidence metric ($\psi$(Xu and Zhang, 2024) to find challenging examples in $\mathcal{D}_{train}$. To be specific, given $x_{i}$ as an input to Grader and $y_{i}$ as its human grading result, we calculate $\psi_{i}=\frac{\max_{\hat{y_{i}}\neq y_{i}}\log{P_{LLM}(\hat{y_{i}}|G,x_{i})}}{\log{P_{LLM}(y_{i}|G,x_{i})}}$, where $\hat{y_{i}}$ is the prediction of Grader. The misconfidence quantifies the discrepancy between the highest log probability of Grader’s incorrect prediction $y_{i}$ and the log probability of correct prediction $\hat{y_{i}}$. Intuitively, the larger $\psi$ indicates that the Grader is giving the wrong judgment with a relatively high confidence over the correct one, thereby implying that the sample is more challenging. However, calculating $\psi$ over all $x_{i}\in\mathcal{D}_{train}$ is computationally expensive and cannot be directly done in each iteration. To avoid introducing the additional computing cost to the current algorithm, we only calculate $\psi_{i}$ for samples in current iteration batch $x_{i}\in b_{out}$ and select the top-$C$ samples as seeds to query similar samples from $\mathcal{D}_{train}$ through embedding similarities. In this way, we simplify the selection process and ensure the consecutive appearance of the similar challenging examples between iterations. At last, to avoid the optimization being operated over the same portion of samples from $\mathcal{D}_{train}$ all the time, we only select half batch based on misconfidence, and keep the another half as random samples. The detailed comparisons between the batch sampling strategies are presented in Section 5.7.
4.3. Test-time Adaptation Stage
@@
In this stage, GradeOpt begins to perform the automatic grading to the large scaled unlabeled responses in test data. However, due to the diversity of language expressions existing in open-ended answers and other influence factors such as geography and time that change users’ expression styles, the performance of the auto-graded is not always guaranteed to be the same as during training. Such phenomenon is well-recognized as the out-of-distribution (OOD) issue in many machine learning problems (Hendrycks and Gimpel, 2016). Prior work (Hendrycks and Gimpel, 2016) has shown that capturing prediction probability statistics about correct or in-sample examples is often sufficient for detecting whether an example is in error or abnormal. Inspired by this, we compose a confidence indicator $\zeta$ as follows: $\zeta=\frac{1}{|\mathcal{D}_{test}|}\sum_{x_{i}\in\mathcal{D}_{test}}\max_{j}(\log P_{LLM}(s_{j}|G,x_{i}))$, where $\log P_{LLM}(\cdot)$ denotes the log likelihood probability given by the LLM. Intuitively, the log probability reflects the confidence that Grader gives to its graded results. By comparing $\zeta$ with the average LLM confidence scores $\mu$ on samples in $\mathcal{D}_{train}$, we can know how serious the OOD phenomenon is. Specifically, when $\zeta>\mu$, it indicates that $G$ is well-applicable to $\mathcal{D}^{\prime}$. When $\zeta<\mu$, it suggests that the guideline is facing serious OOD influences which suggests the current grader may struggle to produce reliable and accurate predictions for $\mathcal{D}_{test}$.
@@
If the test samples are deemed to be OOD, a common solution is to first compose an adaption dataset from the testing scenario. Using this adaption dataset, we then perform test-time training on the existing model. To be specific, test-time training leverages the annotation samples from $\mathcal{D}_{test}$ and fine-tunes the optimized guideline $G_{opt}$ with the same training process introduced in Section 4.2. Unfortunately, in the ASAG scenario, the annotation is usually expensive. Besides, it is challenging to ask pedagogical experts to provide a large amount of annotation samples to help the existing system adapt to any changes in a timely manner. To solve this issue, we propose an incremental labeling approach which checks the marginal performance changes brought by gradually increasing the size of annotation samples. By selecting the size with highest marginal gains in metrics like accuracy and Cohen’s Kappa, GradeOpt only asks pedagogical experts for necessary annotations. This not only reduces the annotation work loads but also increases the adaption efficiency of the framework. Finally, when the $G_{opt}$ passes the OOD test, GradeOpt will be leveraged to finish ASAG over all samples in $\mathcal{D}_{test}$.
5. Experiment

### v2

@@ 4.2.1 Agent Configurations
The Grader focuses on mapping $x_{i}$ to $s_{j}$ based on the given $G$. In GradeOpt, we leverage the exceptional instruction-following capability of LLMs by using a prompt to instruct LLMs to simulate the grading process of human graders. To fully exploit the potential of LLMs, we incorporate the prompt engineering strategy Chain-of-Thought wei2023cot . This encourages LLMs to provide both judgment and intermediate reasoning steps in their outputs. With such design, the Grader becomes better aligned with the human-like grading process. Meanwhile, the intermediate reasoning steps provide support for the Reflector to discover the potential improvements to the given guideline. The prompt for the Grader agent is shown in Figure 3.
Grader Prompt Task Description: In this task, you perform the task of assessing teachers’ knowledge of students’ mathematical thinking by grading teacher’s response to a math teaching question. Question Stem: <question stem> Key Concept: <key concept> Scoring Rubrics: <scoring rubrics> Adaptation Rules: <adaptation rules> Output format <score> Reasoning: <reasoning> Output Rules 1. Replace <score> with only one integer from 0, 1, or 2. 2. Replace <reasoning> with your reasoning. Let’s think step by step!
Figure 3: An example of the prompt to Grader.
@@
The role of Reflector is to propose ways to improve the current guideline $G_{t-1}$ by reflecting over the error samples returned by Grader. To be specific, we design a two-step instruction prompt for LLMs to achieve this goal. In the first step, LLM is instructed to analyze the individual and shared failure reasons for a set of error samples. Then, in the second step, we ask LLMs to propose suggestions that can help resolve those issues. In general, the two-step improving process is analogous to the gradient descent algorithm used by parameter optimization for machine learning algorithm (ruder2017overviewgradientdescentoptimization, ). In our case, the guideline $G$ serves as the parameter of Grader and identifying the error reason is similar to the “gradient". Finally, proposing improving suggestions based on discovered reasons is similar to making a descent down the “gradient" and thus optimizing $G_{t-1}$. The prompt for the Reflector agent is shown in Figure 4.
@@
4.2.2 Iterative Optimization Designs
@@
Reflector Prompt You are ReflectorGPT, a helpful AI agent capable of reflecting on [adaptation rules] that is used by a classifier for a grading task. Your task is to reflect and give reasons for why [adaptation rules] have gotten the given examples in [failed examples] wrong.
The prompt contains two components: 1. [question stem], [key concept] and [scoring rubrics] (these three are given by experts and should not be modified); 2. [adaptation rules] (your task to modify).
Important Steps For Devising Rules: Read [failed examples]. For each one of the errors, perform the following steps: - Step 1: Explain why the classifier made the mistakes, and provide detailed, explanative analyses for why this teacher response should not be interpreted in that wrong way. - Step 2: Devise or modify [adaptation rules] for each mistake to help classifier effectively avoid the mistake and classify the teacher response into the correct category (label). Make sure the devised rule is explanative, straightforward, detailed, concise, and in 1 to 3 sentences.
Question Stem: <question stem> Key Concept: <key concept> Scoring Rubrics: <scoring rubrics> Adaptation Rules: <adaptation rules>
But [adaptation rules] gets the following examples wrong: Failed Examples: <errors>
Give reasons for why [adaptation rules] could have gotten the examples wrong. Let’s think step by step!
Figure 4: An example of the prompt to Reflector.
@@
To accelerate the optimization process and encourage a wider exploration of all possible combinations of error samples in $b_{out}$, we integrate the beam searching strategy (freitag2017beam, ) within both inner and outer loops. The algorithm of the nested iteration is shown in Algorithm 1. To be specific, in the $w$-th inner loop of the $t$-th outer iteration, GradeOpt accepts guidelines beam $G_{t,w-1}=\{g_{t,w-1}^{(k)}\mid 1\leq k\leq K\}$ from $(w-1)$-th inner iteration instead of a single guideline for refining (line 5). Then, during the inner iteration, each $g_{t,w-1}^{(k)}$ will be sent for reflection and refinement with $L$ independently sampled inner batches $b_{in}$ in a parallel manner (line 9). After all refined guidelines for the $w$-th inner loop are finished, $G_{t,w}=\{g_{t,w}^{(l,k)}\mid 1\leq l\leq L,1\leq k\leq K\}$, each new guideline $g_{t,w}^{(l,k)}$ will be tested over a hold-out validation set $\mathcal{D}_{val}$ (line 14). Meanwhile, the top-$K$ performing guidelines will be kept as $G_{t,w}$ and passed to the $(w+1)$-th inner loop. Finally, the beam output of the last iteration of inner loop $G_{t,W}$ will be sent to the $(t+1)$-th outer iteration (line 4).
@@
While this procedure helps increase the accuracy and reliability, blindly increasing the iteration could lead to over-fitting and higher computational overheadsjuneja2024taskfacet . This is particularly true for smaller datasets. To help address these challenges, we introduce an early-stopping criteria. Specifically, during the selection for top-K performed $G_{t,w}$ in the $w$-th inner loop, we record the performance metric $m_{w}$ of the best performed guideline. Then, in the next $(w+1)$-th inner iteration, we check if $m_{w+1}$ is improved. If $m_{w}$ stops improving for two consecutive inner iterations, it indicates that the current guideline is facing risks to be over-fitted, thus following inner iterations are skipped. Similarly, during the $t$-th outer iterations, if its inner iteration is terminated due to the early-stopping and $(t-1)$-th outer iteration’s inner iteration is also terminated by early-stopping, the following outer iterations will also be skipped.
Batch Sampling Strategies
Using self-reflective approaches of LLMs to refine grading guidelines requires the exposure of similar errors in consecutive optimization iterations due to LLMs’ lack of ability in generating appropriate modifications with one attempt ma2024llmsgood . This is especially true for complicated cases involving nuance differences between score categories. However, the randomness of batch sampling in the outer loop fails to guarantee this pre-requisite, which limits the performance of GradeOpt. To solve this, we develop a novel sampling strategy, which leverages the misconfidence metric ($\psi$(xu2024misconfidence, ) to find challenging examples in $\mathcal{D}_{train}$. To be specific, given $x_{i}$ as an input to Grader and $y_{i}$ as its human grading result, we calculate $\psi_{i}=\frac{\max_{\hat{y_{i}}\neq y_{i}}\log{P_{LLM}(\hat{y_{i}}|G,x_{i})}}{\log{P_{LLM}(y_{i}|G,x_{i})}}$, where $\hat{y_{i}}$ is the prediction of Grader. The misconfidence quantifies the discrepancy between the highest log probability of Grader’s incorrect prediction $y_{i}$ and the log probability of correct prediction $\hat{y_{i}}$. Intuitively, the larger $\psi$ indicates that the Grader is giving the wrong judgment with a relatively high confidence over the correct one, thereby implying that the sample is more challenging. However, calculating $\psi$ over all $x_{i}\in\mathcal{D}_{train}$ is computationally expensive and cannot be directly done in each iteration. To avoid introducing the additional computing cost to the current algorithm, we only calculate $\psi_{i}$ for samples in current iteration batch $x_{i}\in b_{out}$ and select the top-$C$ samples as seeds to query similar samples from $\mathcal{D}_{train}$ through embedding similarities. In this way, we simplify the selection process and ensure the consecutive appearance of the similar challenging examples between iterations. At last, to avoid the optimization being operated over the same portion of samples from $\mathcal{D}_{train}$ all the time, we only select half batch based on misconfidence, and keep the another half as random samples. The detailed comparisons between the batch sampling strategies are presented in Section 5.7.
@@
Initialize $G_{0,W}=\{g^{(k)}_{0,W}\}=\{\mathcal{G}\}$;
for $t\leftarrow 1$ to $T$ do
       $b_{out}\leftarrow$ sample an outer iteration batch from $\mathcal{D}_{train}$ ;
       Initialize $G_{t,0}$ = $G_{t-1,W}$ ;
       for $w\leftarrow 1$ to $W$ do
            for $k\leftarrow 1$ to $K$ do
                   $\hat{y}_{out}\leftarrow$ generate grading results for $b_{out}$ by Grader with guideline $g^{(k)}_{t,w}$ ;
                   $e_{t,k}\leftarrow$ find error graded samples from $b_{out}$ caused by guideline $g^{(k)}_{t,w}$ ;
                   do $l\leftarrow 1$ to $L$ in parallel
                         $b_{in}^{(l)}\leftarrow$ randomly sample an inner batch from $e_{t,k}$ ;
                         $g_{t,w}^{(k,l)}\leftarrow$ generate optimized guideline with inputting $b_{in}^{(l)}$ and $g_{t,w-1}^{(k)}$ to Reflector and Refiner ;
            $G_{t,w}=\{g_{t,w}^{(k)}\mid 1\leq k\leq K\}\leftarrow$ select top-$K$ performed $g_{t,w}^{(k,l)}$ based on grading performance over $\mathcal{D}_{val}$ ;
@@
4.3 Test-time Adaptation Stage
@@
In this stage, GradeOpt begins to perform the automatic grading to the large scaled unlabeled responses in test data. However, due to the diversity of language expressions existing in open-ended answers and other influence factors such as geography and time that change users’ expression styles, the performance of the auto-graded is not always guaranteed to be the same as during training. Such phenomenon is well-recognized as the out-of-distribution (OOD) issue in many machine learning problems (hendrycks2016baseline, ). Prior work (hendrycks2016baseline, ) has shown that capturing prediction probability statistics about correct or in-sample examples is often sufficient for detecting whether an example is in error or abnormal. Inspired by this, we compose a confidence indicator $\zeta=\frac{1}{|\mathcal{D}_{test}|}\sum_{x_{i}\in\mathcal{D}_{test}}\max_{j}(\log P_{LLM}(s_{j}|G,x_{i}))$, where $\log P_{LLM}(\cdot)$ denotes the log likelihood probability given by the LLM. Intuitively, the log probability reflects the confidence that Grader gives to its graded results. By comparing $\zeta$ with the average LLM confidence scores $\mu$ on samples in $\mathcal{D}_{train}$, we can know how serious the OOD phenomenon is. Specifically, when $\zeta>\mu$, it indicates that $G$ is well-applicable to $\mathcal{D}^{\prime}$. When $\zeta<\mu$, it suggests that the guideline is facing serious OOD influences, which suggests that grader may struggle to produce reliable and accurate predictions for $\mathcal{D}_{test}$.
@@
If the test samples are deemed to be OOD, a common solution is to first compose an adaption dataset from the testing scenario. Using this adaption dataset, we then perform test-time training on the existing model. To be specific, test-time training leverages the annotation samples from $\mathcal{D}_{test}$ and fine-tunes the optimized guideline $G_{opt}$ with the same training process introduced in Section 4.2. Unfortunately, in the ASAG scenario, the annotation is usually expensive. Besides, it is challenging to ask pedagogical experts to provide a large amount of annotation samples to help the existing system adapt to any changes in a timely manner. To solve this issue, we propose an incremental labeling approach which checks the marginal performance changes brought by gradually increasing the size of annotation samples. By selecting the size with highest marginal gains in metrics like accuracy and Kappa, GradeOpt only asks pedagogical experts for necessary annotations. This not only reduces the annotation work loads but also increases the adaption efficiency of the framework. Finally, when the $G_{opt}$ passes the OOD test, GradeOpt will be leveraged to finish ASAG over all samples in $\mathcal{D}_{test}$.
Refiner Prompt I’m trying to write a classifier for a grading task. You are RefinerGPT, a helpful AI agent capable of refining [adaptation rules] to be used by the classifier.
The [adaptation rules] must contain patterns learned from failed examples, explaining why the predicted score is wrong comparing to the correct label. The set of rules must strictly abide by [scoring rubrics] and must clearly use patterns/details from examples to clearly illustrate and explain. Question Stem: <question stem> Key Concept: <key concept> Scoring Rubrics: <scoring rubrics> Adaptation Rules: <adaptation rules>
But [adaptation rules] have gotten several examples wrong, with the reasons of the problems examined as follows: Failed Examples: <errors> Error Feedbacks: <error feedbacks>
Based on the above information, I wrote one different improved set of rules in replacement of [adaptation rules] for instructing the classifier to learn patterns from examples for avoiding such errors.
Let’s think step by step!
Figure 5: An example of the prompt to Refiner.
5 Experiment